A kite dropped from 168m high cliff and falls with constant velocity of 2.8m/s.32s after kite is relased a pebble is dropped from the cliff. How long does it take for the kite to reach the ground?
1
Expert's answer
2010-08-25T11:33:25-0400
Time needed for the kite to reach the ground: t = s/v = 168/2.8 = 60 (s)
(falling pebble does not affect kite at all)
Answer: 60 s